OpenStudy (anonymous):

does ln e^10=1 or 10?

zepdrix (zepdrix):

Hey you :) Welcome to OpenStudy!

zepdrix (zepdrix):

\[\Large\rm \log(a^b)=b \log(a)\]

zepdrix (zepdrix):

Therefore,\[\Large\rm \ln(e^{10})=10 \ln(e)\]

zepdrix (zepdrix):

Do you remember what happens when the `base` of the log is the same as the inside? :)

zepdrix (zepdrix):

Examples: \[\Large\rm \log_2(2)=1\]\[\Large\rm \log(10)=1\]\[\Large\rm \ln(e)=\log_e(e)=?\]

OpenStudy (anonymous):

well, I know that ln e=1, but depending on the way you enter the equation into the calculator, I get both answers 1 and 10

zepdrix (zepdrix):

This is how you would want to enter it into the calculator: \(\Large\rm \ln(e\text{^}(10))\) Was entering it a different way giving you 1 somehow? :o

OpenStudy (anonymous):

i typed ln (e^10) and got ten, i typed ln (e)^10 and got one, so i wasn't sure which was correct. Thank you :)

OpenStudy (anonymous):

your calculator has two logs one base ten and one base e

zepdrix (zepdrix):

Ahh I see what you did ^^ The second way, your calculator is interpreting it as: \(\Large\rm (\ln e)\text{^}(10)\) So it did the log BEFORE it did the 10, \(\Large\rm (1)\text{^}(10)=1\) Yah that's no bueno c: try to avoid silly mistakes like that hehe.

OpenStudy (anonymous):

we just started this unit today, so i'm a bit new :) i'll try and avoid these mistakes in the future. so to clarify the answer IS 10, correct?

zepdrix (zepdrix):

yes, good c:

OpenStudy (anonymous):

thank you! this helps very much :)